# FlashFormer: Whole-Model Kernels for Efficient Low-Batch Inference

### v1

@@ 4.1 Kernel Benchmark
We apply our kernel to three models: Llama 3.1 8B unquantized, Llama 3.1 8B quantized, and Llama 3.1 70B quantized. We compare with two strong baselines that are widely used in practice: GPTFast [^5] https://github.com/pytorch-labs/gpt-fast and vLLM [^6] https://github.com/vllm-project/vllm [Kwon et al., 2023]. We enable all optimizations where possible. In particular, we enable CUDA graphs to decrease kernel launch overheads in our baselines.
@@
We test a variety of prefill lengths, and generate 128 tokens per output. We measure only the speed of decoding. As a result, we only need to generate enough tokens to achieve a good estimate of the per-token latency with a given amount of prefill. All experiments are performed on one H100 SXM GPU, and use bfloat16 precision [Kalamkar et al., 2019] unless otherwise specified.
Unquantized Llama 3.1 8B: We first evaluate our kernel on an unquantized Llama 3.1 8B model. We find that our kernel achieves a 8-20% speedup over GPTFast, and up to a 61% speedup over vLLM at batch size 1. The results are in Table 5.
Prefill tokens (128 Output) GPTFast vLLM FlashFormer (Ours)
128 168 146 182
1024 159 131 181
3072 158 123 179
6144 143 108 174
Table 5: Tokens ser second (TPS) for Llama 3.1 8B decoding.
@@
Quantized Llama 3.1 8B: We then compare a quantized version of Llama 3.1 8B. Our improvements over baselines in this arena are similar to the unquantized case, ranging from 7-32% as shown in Table 6. We quantize using GPTQ [Frantar et al., 2022] to 4-bit weight only quantization with a group size of 128. While this computation is likely bottlenecked by the operations needed for dequantization and not memory bandwidth, our approach allows us better to overlap dequantization with memory movement and computation.
Prefill tokens (128 Output) GPTFast vLLM FlashFormer (Ours)
128 290 134 311
1024 284 156 310
3072 266 139 307
6144 228 124 301
Table 6: Tokens per second (TPS) for Llama 3.1 8b quantized (4-bit weight-only quantization) decoding.
@@
Quantized Llama 3.1 70B: We finally evaluate on a large quantized model, Llama 3.1 70B. This model is also quantized to 4 bits with GPTQ. Our expectation was that the gains at 70B would be smaller than at 8B. This is because larger models have more opportunities for overlapping within a kernel. This decreases the need for overlapping between kernel launches. The results, shown in Table 7, are small improvements (2%) at short sequence lengths increasing to 13% at longer sequence lengths.
Prefill tokens (128 Output) GPTFast FlashFormer (Ours)
128 47 48
1024 45 48
3072 43 47
6144 40 45
Table 7: Tokens Per Second (TPS) for Llama 3.1 70b quantized decoding. vLLM is omitted due to crashes during quantization on our machine.
4.2 Ablation: Pipeline depth
@@
We investigate how increasing pipeline stages affects kernel performance for the unquantized Llama 3.1 8B model. Our kernel design requires a minimum of two concurrent memory buffers during GLU computation. Consequently, a two-stage pipeline cannot overlap memory operations during GLU in/gate operations. The kernel was primarily optimized for a 64KB pipeline stage size. Performance measurements comparing two versus three pipeline stages at 3072 sequence length reveal that three stages achieves a 5% speedup over two stages, as shown in Table 8.
@@
5 Discussion, & Future Work
5.1 Performance analysis
Two trends emerged in our experiments. First, FlashFormer had the largest speedups at the longest sequence lengths. We suspect this is due to our method of partitioning the KV cache. The size of our pipeline stages is 64KB, which corresponds to 256 keys or values with head dim 128 and bf16 precision. With 8 SMs per KV head, we need 4096 keys and values to provide one matrix chunk per participating SM. At longer sequence lengths this problem vanishes. Instead, FlashFormer overlaps more of the core attention computation and reduction with memory movement, avoiding inefficiency. Future work can explore better optimization of workloads at different sequence lengths.
The second trend we observe is that smaller models benefit more from our technique than larger models. This is likely due to the fact that very large models have more opportunities for overlapping within each operation, and thus gain less benefit from overlapping between operations.
5.2 Comparison to Theoretical Peak
Assuming our model is constrained by memory bandwidth, we can calculate the peak theoretical performance of our model, and what percentage of that performance we observe. If we compute bandwidth utilization as memory movement per token times throughput divided by bandwidth, we calculate that our kernel achieves roughly 82% bandwidth utilization, compared to the 68-75% of GPTFast.
It is unclear what portion of the memory bandwidth is achievable in realistic workloads. A linear layer achieves 86% utilization, while a pure copy kernel achieves  90% bandwidth utilization. As a result, we suspect that our model is starting to approach the practical hardware limits for batch size 1 generation on an H100.
5.3 Compile vs Runtime optimization
@@
However, there are drawbacks to this strategy. At every point of synchronization, the entire GPU is held back by the slowest SM. Furthermore, it can lead to inefficiency when the amount of work changes at runtime, which happens with the KV cache. For our workloads, we tried to avoid this problem through careful division of memory loads. Concurrent work like ThunderMLA [Spector et al., 2023] also implement a dynamic workload partitioning for a persistent FlashAttention style Multi Latent Attention [DeepSeek-AI, 2024] kernel, controlled through a global instruction buffer. Their persistent kernel accesses a global memory buffer to determine what problem to work on next.
@@ 5.4 Limitations
The single-batch setting we target is admittedly limited. We focus on this setting to demonstrate the practicality of the core ideas introduced in this paper, including aggressive specialization and the feasibility of a kernel design with a static workload schedule. We hope that future work applies these ideas to small batch and multi-GPU settings as well as to non-transformer models. Furthermore while we contribute a custom kernel metaprogramming library with customizable kernel templates, further work is needed to extend our abstractions to a greater variety of models and multi-batch / multi-GPU settings.
@@ 6 Related Work
Due to the fact that the linear / GLU layers are standard linear algebra operations, most of the focus of specialized transformer kernel design has focused on optimizing attention. The FlashAttention [Dao et al., 2022, Dao, 2023, Shah et al., 2024b] restructures the attention computation to a numerically equivalent version that fused all of core attention. This general strategy is widespread and was used in many attention or attention-like operations.
Accelerator Kernel DSLs, such as Triton  [Tillet et al., 2019], are used to simplify the kernel generation process by grouping computation into conceptual tiles. Specialized tools such as FlexAttention [Guessous et al., 2024] further allow for simple kernel generation for attention like workloads. Recently, automatic fusion of pointwise operators has continued to improve and become a staple of all modern machine learning libraries. XLA [Sabne, 2020] and PyTorch Compile [Ansel et al., 2024] automate fusion of pointwise operations for Jax [Bradbury et al., 2018] and PyTorch respectively.
Concurrent to this work, ThunderMLA [Spector et al., 2023] introduce a persistent kernel for MultiHead Latent Attention  [DeepSeek-AI, 2024] layers that has each thread block dynamically schedule its own workload.
@@
Speculative decoding [Leviathan et al., 2022] performs autoregressive inference on a small model and multi-token inference on a larger model. This relies on similarity between the model distributions and the relative speed of the small model. The critical idea was that the large verifier model would run at an effective batch size as long as the number of consecutive tokens fed into the model. Conditioned on the tokens verification, a single large model load could generate multiple tokens. Further iterations [Cai et al., 2024, Ankner et al., 2024, Li et al., 2024] improved upon this formula by using draft heads computed on the model’s outputs instead of a separate draft model.
Weight quantization methods [Frantar et al., 2022, Yao et al., 2022, Chee et al., 2023] have accelerated decoding by enabling lower and lower bit precisions for weight loads. This decreases the cost of memory movement, and thus improves inference speed. However, this is not perfect, as most of these methods utilize scarce GPU resources for dequantization.
@@ 7 Conclusion
Acknowledgments
This study was supported by MIT-IBM Watson AI Lab.
Appendix A Detailed Algorithm Diagrams
Figure 5: The Attention forward pass per layer. Large matrices (green) are loaded asynchronously from global, and do not wait on consumers. The consumers works synchronously, with 3 global synchronizations per attention forward pass. Each of the repeated computations (in red) corresponds to a asynchronously loaded matrix split into chunks.

### v2

@@ 5.1 Experimental Setup
We apply our kernel to three models: Llama 3.1 8B unquantized, Llama 3.1 8B quantized, and Llama 3.1 70B quantized. We compare with two strong baselines that are widely used in practice: GPTFast [^4] https://github.com/pytorch-labs/gpt-fast and vLLM [^5] https://github.com/vllm-project/vllm (Kwon et al., 2023). We enable all optimizations where possible. In particular, we enable CUDA graphs to decrease kernel launch overheads in our baselines.
@@
We test a variety of prefill lengths, and generate 128 tokens per output. We only measure decoding speed, as prefill is compute bound. As a result, we only need to generate enough tokens to achieve a good estimate of the per-token latency with a given amount of prefill. All experiments are performed on one H100 SXM GPU.
5.2 Main Results
Batch size = 1 case.
@@ 5.2 Main Results
We first focus on the batch size = 1 setting and show the results in Table˜6 and Figure˜7. On unquantized Llama 3.1 8B, we find that our kernel achieves a 8-20% speedup over GPTFast, and up to a 61% speedup over vLLM. We observe similar improvements with quantization, where we use GPTQ (Frantar et al., 2022) to perform weight-only integer quantization to 4 bits with a group size of 128. While this computation is likely bottlenecked by the operations needed for dequantization and not memory bandwidth, our approach allows us better to overlap dequantization with memory movement and computation.
@@
We also evaluate on a large quantized model, Llama 3.1 70B. Our expectation was that the gains at 70B would be smaller than at 8B. This is because larger models have more opportunities for overlapping within a kernel. This decreases the need for overlapping between kernel launches. The results, shown in Table˜6, are small improvements (2%) at short sequence lengths increasing to 13% at longer sequence lengths.
Larger batch sizes.
We perform experiments with batch size two and four with the unquantized 8B model. We consider the case where KV caches are entirely distinct between elements of the batch, and generation for all queries within the batch starts and stops at the same time. These batched kernels are still input stationary: weights are read in exactly once, and inputs are fully replicated on each SM. The synchronization and bandwidth savings from the batch size one case hold in the slightly larger batch setting, as shown in Table˜7. However, memory bandwidth utilization decreases, as the increased shared memory pressure does force smaller memory pipelines or a minimal L1 cache.
Discussion.
Two trends emerged in our experiments. First, FlashFormer had the largest speedups at the longest sequence lengths. We suspect this is due to our method of partitioning the KV cache. The size of our pipeline stages is 64KB, which corresponds to 256 keys or values with head dim 128 and BF16 precision. With 8 SMs per KV head, we need 4096 keys and values to provide one matrix chunk per participating SM. At longer sequence lengths this problem vanishes. Instead, FlashFormer overlaps more of the core attention computation and reduction with memory movement, improving bandwidth utilization. Future work can explore better optimization of workloads at different sequence lengths.
The second trend we observe is that smaller models benefit more from our technique than larger models, as evidenced by Figure˜7. This is likely due to the fact that very large models have more opportunities for overlapping within each operation, and thus gain less benefit from overlapping between operations.
Batch Size SL GPTFast vLLM FlashFormer
2 128 308 286 328
4 128 608 564 616
2 1024 556 560 648
4 1024 504 548 600
Table 7: Batched tokens per second for unquantized Llama 3.1 8B under various settings. SL Stands for sequence length.
5.3 Ablations & Analysis
We next perform a series of ablations and analyses.
Pipeline depth.
@@ 5.3 Ablations & Analysis
We investigate how increasing pipeline stages affects kernel performance for the unquantized Llama 3.1 8B model. Our kernel design requires a minimum of two concurrent memory buffers during GLU computation. Consequently, a two-stage pipeline cannot overlap memory operations during GLU in/gate operations. The kernel was primarily optimized for a 64KB pipeline stage size. Performance measurements comparing two versus three pipeline stages at 3072 sequence length reveal that three stages achieves a 5% speedup over two stages, as shown in Table˜8.
@@
Attention & GLU kernels.
We measure the latency and memory bandwidth of stacked layers of Attention and GLU blocks, similar to our motivating study with stacked linear layers. The results are in Table˜9 and Figure˜8.
For latency, we see that double digit latency improvements occur from cross layer fusion, and single digit improvements occur from cross layer memory overlapping, regardless of layer type. Additionally, we ablate the width and depth of our pipeline for the Attention and GLU kernels. While the the attention and linear layers are fastest with a depth 2 pipeline, the GLU layer is fastest with a depth 3 pipeline. As GLU represents the majority of parameters and latency, the composite kernel is fastest with a depth 3 pipeline. Future work may look at adaptively changing pipeline size throughout the model
For memory bandwidth, while stacked GLU blocks are able exceed 90% of machine peak, our stacked attention blocks achieve roughly 50% of machine peak. Compared to GLU layers, attention layers have more logic, special functions, and synchronization with less memory movement to hide them. Note that as nearly eighty percent of layer parameters are in GLU layers, the bandwidth of our resulting kernel is much closer to the bandwidth of the stacked GLU layer.
Module Layers Baseline Fused +Overlap
1 0.089 0.089 0.090
Attn 4 0.357 0.264 0.256
32 2.857 1.761 1.683
1 0.122 0.122 0.122
GLU 4 0.488 0.466 0.464
32 3.905 3.689 3.661
Table 9: Component-wise latencies for multilayer attention and GLU kernels, in ms. Baseline does not fuse any layers. Fused refers to a kernel that fuses across layers without memory overlapping. + Overlap incorporates cross-layer memory overlapping.
6 Discussion & Limitations
6.1 Handwritten vs Compiled Kernels
One interesting direction of research is automatically compiling whole-model kernels. Mirage (Wu et al., 2025) recently added support to user their kernel generator to automatically generate whole model kernels. We implemented an optimized Llama3 in for Hopper GPUs in their framework and compared TPS numbers. Unfortunately their benchmarking tools combine prefill and decoding latency, but even when generating without any prefill the peak performance we achieved was 153 TPS with LLama3 8B. The difference in performance appears to be due to better work partitioning, decreased synchronization, and greater overlapping in our kernel. Currently, the greater exploration space opened up by handwriting kernels leads to  20% performance gains in our case.
Figure 8: Attention and GLU memory bandwidth in an unquantized Llama 3.1 8B kernel. While GLU nears hardware peak, our attention implementation achieves closer to 50% of peak bandwidth. As GLU dominates memory usage, net bandwidth is much closer to GLU memory bandwidth.
6.2 Compile vs. Runtime optimization
@@ 6.1 Handwritten vs Compiled Kernels
However, there are drawbacks to this strategy. At every point of synchronization, the entire GPU is held back by the slowest SM. Furthermore, it can lead to inefficiency when the amount of work changes at runtime, which happens with the KV cache. For our workloads, we tried to avoid this problem through careful division of memory loads. Works such as ThunderMLA (Spector et al., 2023) also implement a dynamic workload partitioning for a persistent FlashAttention style Multi Latent Attention (DeepSeek-AI, 2024) kernel, controlled through a global instruction buffer. Their persistent kernel accesses a global memory buffer to determine what problem to work on next.
@@ 6.3 Limitations
Our work is limited insofar as we target the single-GPU, low-batch setting for the Hopper family of architectures. We focus on this setting to demonstrate the practicality of the core ideas introduced in this paper, including aggressive specialization and the feasibility of a kernel design with a static workload schedule. We hope that future work applies these ideas to larger-batch, multi-GPU settings on different families or architectures, as well as to non-transformer models. Furthermore while we contribute a custom kernel metaprogramming library with customizable kernel templates, further work is needed to extend our abstractions to a greater variety of models and multi-GPU settings.
@@ 7 Related Work
Due to the fact that the linear / GLU layers are standard linear algebra operations, most of the focus of specialized transformer kernel design has focused on optimizing attention. The FlashAttention (Dao et al., 2022; Dao, 2023; Shah et al., 2024b) restructures the attention computation to a numerically equivalent version that fused all of core attention. This general strategy is widespread and was used in many attention or attention-like operations.
Accelerator Kernel DSLs, such as Triton  (Tillet et al., 2019), are used to simplify the kernel generation process by grouping computation into conceptual tiles. Specialized tools such as FlexAttention (Guessous et al., 2024) further allow for simple kernel generation for attention like workloads. Recently, automatic fusion of pointwise operators has continued to improve and become a staple of all modern machine learning libraries. XLA (Sabne, 2020) and PyTorch Compile (Ansel et al., 2024) automate fusion of pointwise operations for Jax (Bradbury et al., 2018) and PyTorch respectively.
Concurrent to this work, ThunderMLA (Spector et al., 2023) introduce a persistent kernel for MultiHead Latent Attention  (DeepSeek-AI, 2024) layers that has each thread block dynamically schedule its own workload.
@@
Speculative decoding (Leviathan et al., 2022) performs autoregressive inference on a small model and multi-token inference on a larger model. This relies on similarity between the model distributions and the relative speed of the small model. The critical idea was that the large verifier model would run at an effective batch size as long as the number of consecutive tokens fed into the model. Conditioned on the tokens verification, a single large model load could generate multiple tokens. Further iterations (Cai et al., 2024; Ankner et al., 2024; Li et al., 2024) improved upon this formula by using draft heads computed on the model’s outputs instead of a separate draft model.
Weight quantization methods (Frantar et al., 2022; Yao et al., 2022; Chee et al., 2023) have accelerated decoding by enabling lower and lower bit precisions for weight loads. This decreases the cost of memory movement, and thus improves inference speed. However, this is not perfect, as most of these methods utilize scarce GPU resources for dequantization.
@@ 8 Conclusion
Acknowledgements
This work was supported by the National Science Foundation under CAREER Award No. 2441872 and the MIT-IBM Watson AI Lab.
Appendix A Metaprogramming Language Example
LABEL:tab:cheetah_advanced shows more examples from our Cheetah metaprogramming language.
Python Implementation Generated CUDA Kernel
class AsyncBarrierGTSCopyFn(SmartScopeFn): ... def setup( self, src_cfg: SDIPConfig, dst_cfg: SDIPConfig, thread_idx: DimName, cpy_idx: DimName, ): ... def generate( self, ix: Indices, src_ptr: SafeDataIndexPtr, dst_ptr: SafeDataIndexPtr, barrier: SharedMBarrier, ): size = (ix.size(self.cpy_idx) * dst_ptr.get_elem_size()) ix.set_index(self.cpy_idx, ch.const(0, ty.i32)) assert size % 16 == 0 with ch.if_(ix[self.thread_idx] == 0): ch.asm( f"cp.async.bulk.shared::cluster.global.mbarrier::complete_tx::bytes [\$dst_ptr], [\$src_ptr], {size}, [\$barrier];", dst_ptr=dst_ptr.idx_offset(), src_ptr=src_ptr.idx_offset(), barrier=barrier.ptr, ) ch.asm( f"mbarrier.expect_tx.relaxed.cta.shared.b64 [\$barrier], {size};", barrier=barrier.ptr, ) ch.raw_stmt("__syncwarp();") ... async_gts_copy_fn(ix, src_ptr, shared_ptr, shared_barrier) int32_t elem_idx0 = 0; if (thread_idx == 0) { asm ("cp.async.bulk.shared::cluster.global.mbarrier::complete_tx::bytes [%0], [%1], 32768, [%2];" : : "l"(&shared[elem_idx0]), "l"(&x[elem_idx0]), "r"(mbar_shmem)); asm ("mbarrier.expect_tx.relaxed.cta.shared.b64 [%0], 32768;" : : "r"(mbar_shmem)); } __syncwarp(); self.qk_compute_fn( ix, self.shared_query_buffer_ptr, k_buffer, self.scores_ptr, self.mem_pipeline.consumer_barrier, )